find the inverse to the matrix

1 -1 2
N= -1 3 2
2 1 -3

and use this inverse to solve the system of equations
x-y+2z=1
-x+3y+2z=1
2x+y-3z=5
1
Expert's answer
2012-10-11T09:26:25-0400
The inverse matrix is

11/26 1/26 4/13
N^(-1) = -1/26 7/26 2/13
7/26 3/26 -1/13


By the way, you can watch our video explaining how to find inverse matrices



Let's solve the given system:

(1)
(x, y, z) = (1) * N^(-1) = (45/26, 23/26, 1/13).
(5)